Electromagnetic radiation is emitted when a neutron and proton fuse to make a deuteron. What is the wavelength of this radiation? Number 5.136 x 10 nm
Answer #1

mass of proton = 1.6726 x 10-27 kg

mass of neutron = 1.6749 x 10-27 kg

mass of deuteron= 3.34 x 10-27 kg

Missing mass = (1.6726 +1.6749 - 3.34) x 10-27 kg = 0.0075 x 10-27 kg


Energy released E = mc2 , c is velocity of light = 3 x 108 m/s

                              = 0.0075 x 10-27 kg x (3 x 108 m/s)2

                               = 6.75 x 10-13 J


We know that E = hc/wave length   , h = planck’s constant = 6.626 x 10-34 J.s

Then,
wave length = hc/E

                      = [6.626 x 10-34 J.s x 3x 108 m/s] /6.75 x 10-13 J

                      = 2.94 x 10-13 m

                  = 2.94 x 10-4 x 10-9 m

                      = 2.94 x 10-4 nm                                 [ 1 nm = 10-9 nm]

Therefore,

wave length of given radiation = 2.94 x 10-4 nm